# Transformation of orientation and rotation angles of synchronous satellites: Application to the Galilean moons

### v1

@@ 7 Recommendations for updating the IAU WGCCRE model
The current IAU WG solution includes a limited number of terms and is provided with a presumed accuracy of $0.01^{\circ}$ on each nutation amplitudes (we assumed the precision is one order higher than the given number of digits). Our solution is by construction more accurate (to within about $0.001^{\circ}$ degrees for the time-series in $\alpha$, due to the truncation error of Table 9), but is based on the assumption that the satellites are entirely rigid. Therefore, it is not certain that, for the time being, an update of the IAU WG solution should include as many terms as ours. If a simplified model with an accuracy of the order of $0.001^{\circ}$ on each nutation term is considered sufficient for the four satellites, some simplifications can be made to the periodic series of Table 9. Firstly, we propose to retain only the $5$, 99, 77, and $4$ largest terms, respectively, for the four satellites. Thus, the $5.93$-year periodicity reported for Callisto in Archinal et al. [2018] and also observed in Table 4 (third line) is neglected here (its period is much shorter than the free precession period of approximately $204$ years, so that the associated inertial obliquity amplitude is significantly smaller than the inclination amplitude). Next, we suggest reducing the number of arguments used to express the solution. Remember that some arguments listed in Table 9 correspond to identical combinations of fundamental frequencies but may differ slightly in value between satellites due to fitting to numerical integrations (e.g., $\dot{\Omega}_{4}=-0.01113$ rad/year in the Io series, versus 0.01120-0.01120 rad/year in the Callisto series). When such differences do not degrade the model’s accuracy beyond the $0.0005$-degree threshold, we use a common value for the argument across satellites. For example, for $\Omega_{1}$, which dominates the periodic series of Io, but is only fifth in the series for Europa, we choose the value of the argument as found Io’s series, see Table 12. The differences from the original arguments are usually smaller than $10^{-4}$ rad/year and $0.5^{\circ}$ in frequency and phase, respectively. Only differences in phase value can cause differences in the time domain that exceed the threshold. The precision of the time series obtained using $14$ different arguments rather than the $25$ arguments of Table 9 is smaller than $0.0002^{\circ}$. The proposed update for the IAU WG solution, with the simplified periodic series, is given in Table 12. There is no physical libration in the expression for $W$, only periodic variations due to the projection of the spin precession/nutations. The differences between our model based on the JUP orbital theory and the IAU WG solution of Archinal et al. [2018] can reach up to $0.4^{\circ}$ after $100$ years, see Fig. 9.
@@
Figure 9: Differences in right ascension, declination, and rotation angle between our model and the IAU WG solution [Archinal et al., 2018] for each of the four Galilean satellites.
@@ 8.1 Sensitivity to the choice of ephemeris
$1/\cos\delta_{LP}$ 2.3226 2.3236 2.3285 2.3504
$\tan\delta_{LP}\,/\cos\delta_{LP}$ 4.8687 4.8736 4.8965 4.9995
$1/2\,\tan\delta_{LP} -$1.0481 -1.0487 -1.0514 -1.0635
$\cos\delta_{LP}$ 0.4306 0.4304 0.4295 0.4255
$\sin\delta_{LP} -$0.9026 -0.9027 -0.9031 -0.9050
$1/4\sin 2\delta_{LP}$ 0.1943 0.1942 0.1939 0.1925
@@
A comparison between the orbit orientations of the Galilean satellites obtained with JUP387 or NOE reveals differences, on both short and long time scales, see Fig. 10. As with JUP387, we numerically fit the Laplace plane coordinates ($\alpha_{LP},\delta_{LP}$), a periodic series and one secular term for the projected coordinates ($n_{x},n_{y}$) of the orbit normal onto the Laplace plane. The Laplace plane coordinates differ by as much as $0.004^{\circ}$ from those associated to JUP387, which also affects some of the transformation coefficients, see Table 13. The small secular rates ($tr_{x/y}$) associated to the two ephemerides differ significantly, see Tables 3 and 15, with discrepancies in $n_{x/y}$ reaching up to $0.005^{\circ}$ after $100$ years, corresponding to a difference up to $150$ km in the position of the Galilean satellite around Jupiter. The differences in the amplitudes of the $n_{x/y}$ series between JUP387 and NOE range from approximately $3\,10^{-4}$ deg (for Io) to $2\,10^{-3}$ deg (for Callisto), see Tables 4 and 14. Note that for each satellite, since the Laplace planes of JUP387 and NOE are different, the $n_{x/y}$ coordinates corresponding to each ephemerides are expressed in a different reference frame. The transformation into equatorial coordinates $\alpha_{O}$ and $\delta_{O}$ allows for a comparison of the orbit orientation in an identical reference frame (see gray and orange curves in Fig. 10).
@@
Io $7.3\,10^{-5}$ $2.8\,10^{-6}$ $1.7\,10^{-4}$ $1.5\,10^{-4}$ 203.4889584317 $1.4\,10^{-4}$
Europa $7.3\,10^{-5}$ $3.0\,10^{-6}$ $1.7\,10^{-4}$ $1.5\,10^{-4}$ 101.3747246256 $2.5\,10^{-3}$
Ganymede $7.2\,10^{-5}$ $4.1\,10^{-6}$ $1.7\,10^{-4}$ $1.5\,10^{-4}$ 50.3176077039 $4.6\,10^{-4}$
Callisto $6.7\,10^{-5}$ $8.6\,10^{-6}$ $1.6\,10^{-4}$ $1.4\,10^{-4}$ 21.5710726895 $1.4\,10^{-3}$
Figure 11: Comparison of the orbit and the spin position (for a solid satellite and one with a liquid layer) in the right ascension/declination plane for Europa. Three orbital theories are compared: JUP387 (blue), NOE (red, below the JUP387 curve), IAU WG model (cyan). The spin position is plotted in a solid case (brown) and a liquid case (orange, obliquity of 0.1 deg). The time range covered by the plot is 2015.5-2023.5. The black dots are evaluated in 2020.0. The instantaneous Cassini plane is also plotted with dotted lines. Different Laplace poles are also shown.
@@
Figure 11 shows the orbit normal and spin motion of Europa in the ICRF plane over a $9$-year period. The angular separation between the spin axis and the orbit normal (also referred to as the orbital obliquity) is clearly noticeable. The spin motion is shown for both a solid interior model and an interior model with an internal global liquid ocean (see discussion in Section 8.2). The difference between the JUP and NOE models is not noticeable at this scale. As expected, the IAU WG model remains close to the orbit normal path, but deviates significantly from the spin axis trajectory. For reference, the solutions of Bills and Scott [2022] for the orbit normal and spin motion, which are inconsistent with both the IAU WG model and our solutions, are also shown in the figure.
@@ 8.2 Effect of the interior model
Table 18: Amplitudes $\varepsilon_{j}$, $\alpha_{S}$ and $\delta_{S}$ for Europa, Ganymede and Callisto for a range of interior models with a liquid layer for the JUP387 orbital theory, after removing the most extreme $5\%$ of values from each range. Only ranges whose limits are relatively far from the solid value and of the order of at least about $0.01$ degrees are kept here. The ranges in $\mu_{j}$ can be approximately estimated by scaling the $\alpha^{S}_{j}$ column by a factor of 0.9-0.9. For Europa and Callisto, the ranges in corrective term $\delta_{S\,0}$ are also affected by the resonances and are equal to the ranges in the amplitudes $\delta^{S}_{4}$ and $\delta^{S}_{3}$, respectively, of the term with twice the main frequency (order 2 term).
@@
We consider a wide range of interior structure models for Europa, Ganymede and Callisto, consistent with known constraints (mass, radius, and MOI), by varying the densities and sizes of their different layers in ranges similar to those of Baland et al. [2012] or Coyette et al. [2025] and use the theoretical angular momentum model of Baland et al. [2019] for a synchronous satellite with an internal global fluid layer to compute the ranges of obliquity amplitudes and build the series for $s_{x/y}$. The series for $\alpha,\delta,$ and $\mu$ are then obtained with the transformations described in Sections 3 and 4. As anticipated in Section 6.2, the amplitudes of the spin axis motion at a given forcing frequency can differ significantly from the values obtained in the solid case when one of the free frequencies associated with the interior is sufficiently close to the forcing frequency. In such cases, the ordering of the terms may differ from the ordering in the solid case, and the approximation for $\varepsilon(t),\theta(t),\zeta(t)$, and $\psi(t)$ that assume the existence of a dominant amplitude (e.g. Eqs. 51) become ineffective when two amplitudes dominate the solution together. Because of resonant amplification, obliquity amplitudes can reach very large absolute values, which can be considered unrealistic given that tidal damping is not taken into account in the dynamic model. Therefore, we remove the most extreme $5\%$ of values from each range of obliquity amplitudes before calculating the minimum and maximal values from the remaining set. This gives us indicative limits for the obliquity, right ascension, and declination amplitudes, see Table 18, which do not necessarily imply that an amplitude cannot be larger or smaller than these limits.
@@ 9 Conclusions and discussion
A solution similar to that of Archinal et al. [2018] can be recovered by reducing the number of frequencies and assuming certain close phases and frequencies to be equal. Our proposed model now includes 14 frequencies shared among the four Galilean satellites, compared to only 6 in the IAU WG model (Section 7). In our solution, frequencies have a sign assigned based on the original frequency in the forcing series (corresponding to prograde or retrograde motions), which is crucial for comparison with the free retrograde mode and for evaluating possible amplifications. The discrepancies between our model and the Lieske [1979] model incorporated in Archinal et al. [2018] can reach up to $0.4^{\circ}$ after $100$ years. These differences arise from the use of an updated orbital theory and an non-zero obliquity model (whereas the previous IAU WG solution assumed zero obliquity, our model incorporates a theoretical obliquity consistent with a solid, rigid satellite). As in Archinal et al. [2018], our model contains no forced librations in the $W$ series; the periodic variations in $W$ originate solely from the periodic terms of the $\mu$ angle.
Our nominal series are based on JUP387, but we also provide equivalent series derived from the NOE ephemerides (Section 8.1). Although the frequencies, phases, and amplitudes differ slightly between the two series, the largest deviations in the spin position are due to the trend differences in the orbital precession. A SPICE kernel for the case of entirely solid satellites has been generated for the $\alpha,\delta$ and $W$ angles, using both the JUP and NOE ephemerides. These 2 kernels are currently available on lara.oma.be/GalileanSat/. Another localisation on the SPICE kernel architecture will be available soon. If Europa, Ganymede, or Callisto possess a liquid global ocean beneath the icy shell, certain amplitudes may be amplified, as summarized in Table 18.
@@ Acknowledgment
We thank Sébastien Le Maistre, Tim Van Hoolst and Valery Lainey for their helpful comments and discussions. We thank Alfonso Caldiero for testing the SPICE kernel.

### v2

@@ 7 Recommendations for updating the IAU WGCCRE solution
The current IAU WG solution includes a limited number of terms and is provided with a presumed accuracy of $0.01^{\circ}$ on each nutation amplitudes (we assumed the precision is one order higher than the given number of digits). Our solution is by construction more accurate (to within about $0.001^{\circ}$ degrees for the time-series in $\alpha$, due to the truncation error of Table 9), but is based on the assumption that the satellites are entirely rigid. Therefore, it is not certain that, for the time being, an update of the IAU WG solution should include as many terms as ours. If a simplified model with an accuracy of the order of $0.001^{\circ}$ on each nutation term is considered sufficient for the four satellites, some simplifications can be made to the periodic series of Table 9. Firstly, we propose to retain only the $5$, 99, 77, and $4$ largest terms, respectively, for the four satellites. Thus, the $5.93$-year periodicity reported for Callisto in Archinal et al. [2018] and also observed in Table 4 (third line) is neglected here (its period is much shorter than the free precession period of approximately $204$ years, so that the associated inertial obliquity amplitude is significantly smaller than the inclination amplitude). Next, we suggest reducing the number of arguments used to express the solution. Remember that some arguments listed in Table 9 correspond to identical combinations of fundamental frequencies but may differ slightly in value between satellites due to fitting to numerical integrations (e.g., $\dot{\Omega}_{4}=-0.01113$ rad/year in the Io series, versus 0.01120-0.01120 rad/year in the Callisto series). When such differences do not degrade the model’s accuracy beyond the $0.0005^{\circ}$ threshold, we use a common value for the argument across satellites. For example, for $\Omega_{1}$, which dominates the periodic series of Io, but is only fifth in the series for Europa, we choose the value of the argument as found Io’s series, see Table 12. The differences from the original arguments are usually smaller than $10^{-4}$ rad/year and $0.5^{\circ}$ in frequency and phase, respectively. Only differences in phase value can cause differences in the time domain that exceed the threshold. The precision of the time series obtained using $14$ different arguments rather than the $25$ arguments of Table 9 is smaller than $0.0002^{\circ}$. The proposed update for the IAU WG solution, with the simplified periodic series, is given in Table 12. There is no physical libration in the expression for $W$, only periodic variations due to the projection of the spin precession/nutations. The differences between our model based on the JUP orbital theory and the IAU WG solution of Archinal et al. [2018] can reach up to $0.4^{\circ}$ after $100$ years, see Fig. 10.
@@
Figure 10: Differences in right ascension, declination, and rotation angle between our model and the IAU WG solution [Archinal et al., 2018] for each of the four Galilean satellites. The purple box represents the timing of the JUICE mission.
@@ 8.1 Sensitivity to the choice of ephemeris
$1/\cos\delta_{LP}$ $\mathcal{O}1$ 2.3226 2.3236 2.3285 2.3504
$\tan\delta_{LP}\,/\cos\delta_{LP}$ $\mathcal{O}2$ 4.8687 4.8736 4.8965 4.9995
$1/2\,\tan\delta_{LP}$ $\mathcal{O}2$ -1.0481 -1.0487 -1.0514 -1.0635
$\cos\delta_{LP}$ $\mathcal{O}1$ 0.4306 0.4304 0.4295 0.4255
$\sin\delta_{LP}$ $\mathcal{O}2$ -0.9026 -0.9027 -0.9031 -0.9050
$1/4\sin 2\delta_{LP}$ $\mathcal{O}2$ 0.1943 0.1942 0.1939 0.1925
@@
Io $7.3\times 10^{-5}$ $2.8\times 10^{-6}$ $1.7\times 10^{-4}$ $1.5\times 10^{-4}$ 203.4889584317 $1.4\times 10^{-4}$
Europa $7.3\times 10^{-5}$ $3.0\times 10^{-6}$ $1.7\times 10^{-4}$ $1.5\times 10^{-4}$ 101.3747246256 $2.5\times 10^{-3}$
Ganymede $7.2\times 10^{-5}$ $4.1\times 10^{-6}$ $1.7\times 10^{-4}$ $1.5\times 10^{-4}$ 50.3176077039 $4.6\times 10^{-4}$
Callisto $6.7\times 10^{-5}$ $8.6\times 10^{-6}$ $1.6\times 10^{-4}$ $1.4\times 10^{-4}$ 21.5710726895 $1.4\times 10^{-3}$
A comparison between the orbit orientations of the Galilean satellites obtained with JUP387 or NOE reveals differences, on both short and long time scales, see Fig. 11. As with JUP387, we numerically fit the Laplace plane coordinates ($\alpha_{LP},\delta_{LP}$), a periodic series and one secular term for the projected coordinates ($n_{x},n_{y}$) of the orbit normal onto the Laplace plane. The Laplace plane coordinates differ by as much as $0.004^{\circ}$ from those associated to JUP387, which also affects some of the transformation coefficients, see Table 13. The small secular rates ($tr_{x/y}$) associated to the two ephemerides differ significantly, see Tables 3 and 14, with discrepancies in $n_{x/y}$ reaching up to $0.005^{\circ}$ after $100$ years, corresponding to a difference up to $150$ km in the position of the Galilean satellite around Jupiter. The differences in the amplitudes of the $n_{x/y}$ series between JUP387 and NOE range from approximately $3\times 10^{-4}$ deg (for Io) to $2\times 10^{-3}$ deg (for Callisto), see Tables 4 and 15. The differences in the periods are the largest for term $j=5$ of Io’s series ($7$ years of difference). This term, with argument $\Omega_{4}$, has a small amplitude and a long period and is therefore more difficult to estimate uniquely than a term with a large amplitude and short period. For the other satellites, the $\Omega_{4}$ terms are relatively more important, and their period is generally better constrained to around 560-560 years. Note that for each satellite, since the Laplace planes of JUP387 and NOE are different, the $n_{x/y}$ coordinates corresponding to each ephemerides are expressed in a different reference frame. The transformation into equatorial coordinates $\alpha_{O}$ and $\delta_{O}$ allows for a comparison of the orbit orientation in an identical reference frame (see gray and orange curves in Fig. 11).
@@
Figure 12: Comparison of the orbit and the spin position (for a solid satellite and one with a liquid layer) in the right ascension/declination plane for Europa. Three orbital theories are compared: JUP387 (blue), NOE (red, below the JUP387 curve), IAU WG solution (cyan). The spin position is plotted in a solid case (brown) and a liquid case (green, obliquity of 0.1 deg). The time range covered by the plot is 2011-2025. The black dots are evaluated in 2020.0. The instantaneous Cassini plane is also plotted with dotted lines. The Laplace plane associated with each orbit is given in the same color as the orbit.
@@
Figure 12 shows the orbit normal and spin motion of Europa in the ICRF plane over a $9$-year period. The angular separation between the spin axis and the orbit normal (also referred to as the orbital obliquity) is clearly noticeable. The spin motion is shown for both a solid interior model and an interior model with an internal global liquid ocean (see discussion in Section 8.2). The difference between the JUP and NOE models is not noticeable at this scale. As expected, the IAU WG model remains close to the orbit normal path, but deviates significantly from the spin axis trajectory. For reference, the solutions of Bills and Scott [2022] for the orbit normal and spin motion, which are inconsistent with both the IAU WG solution and our solutions, are also shown in the figure.
@@ 8.2 Effect of the interior model
Table 18: Amplitudes $\varepsilon_{j}$, $\alpha_{S}$ and $\delta_{S}$ for Europa, Ganymede and Callisto for a range of interior models with a liquid layer for the JUP387 orbital theory, after removing the most extreme $5\%$ of values from each range. Only ranges whose limits are clearly far from the solid value and of the order of at least about $0.01$ degrees are kept here. The ranges in $\mu_{j}$ can be approximately estimated by scaling the $\alpha^{S}_{j}$ column by a factor of 0.9-0.9. For Europa and Callisto, the ranges in corrective term $\delta_{S\,0}$ are also affected by the resonances and are equal to the ranges in the amplitudes $\delta^{S}_{4}$ and $\delta^{S}_{3}$, respectively, of the term with twice the main frequency (order 2 term).
@@
We consider a wide range of interior structure models for Europa, Ganymede and Callisto, consistent with known constraints (mass, radius, and MOI), by varying the densities and sizes of their different layers in ranges similar to those of Baland et al. [2012] or Coyette et al. [2026] and use the theoretical angular momentum model of Baland et al. [2019] for a synchronous satellite with an internal global fluid layer to compute the ranges of obliquity amplitudes and build the series for $s_{x/y}$. The series for $\alpha,\delta,$ and $\mu$ are then obtained with the transformations described in Sections 3 and 4. As anticipated in Section 6.2, the amplitudes of the spin axis motion at a given forcing frequency can differ significantly from the values obtained in the solid case when one of the free frequencies associated with the interior is sufficiently close to the forcing frequency. In such cases, the ordering of the terms may differ from the ordering in the solid case, and the approximation for $\varepsilon(t),\theta(t),\zeta(t)$, and $\psi(t)$ that assume the existence of a dominant amplitude (e.g. Eqs. 52) become ineffective when two amplitudes dominate the solution together. Because of resonant amplification, obliquity amplitudes can reach very large absolute values, which can be considered unrealistic given that tidal damping is not taken into account in the dynamic model. Therefore, we remove the most extreme $5\%$ of values from each range of obliquity amplitudes before calculating the minimum and maximal values from the remaining set. The quoted values represent a typical 95% interval rather than strict limits, giving us an indicative range for the obliquity, right ascension, and declination amplitudes, see Table 18, which do not necessarily imply that an amplitude cannot be larger or smaller than these limits.
@@ 9 Conclusions and discussion
A solution similar to that of Archinal et al. [2018] can be recovered by reducing the number of frequencies and assuming certain close phases and frequencies to be equal. Our proposed model now includes 14 frequencies shared among the four Galilean satellites, compared to only 6 in the IAU WG solution (Section 7). In our solution, frequencies have a sign assigned based on the original frequency in the forcing series (corresponding to prograde or retrograde motions), which is crucial for comparison with the free retrograde mode and for evaluating possible amplifications. The discrepancies between our model and the Lieske [1979] model incorporated in Archinal et al. [2018] can reach up to $0.4^{\circ}$ after $100$ years. These differences arise from the use of an updated orbital theory and an non-zero obliquity model (whereas the previous IAU WG solution assumed zero obliquity, our model incorporates a theoretical obliquity consistent with a solid, rigid satellite). As in Archinal et al. [2018], our model contains no forced librations in the $W$ series; the periodic variations in $W$ originate solely from the periodic terms of the $\mu$ angle. Because the series in $\alpha$, $\delta$, and $\mu$ have mostly long term periodicities, we do not expect them to be well constrained by a single spacecraft mission, which is typically of limited duration. In such cases, a local approximation is likely more appropriate.
Our nominal series are based on JUP387, but we also provide equivalent series derived from the NOE ephemerides (Section 8.1). Although the frequencies, phases, and amplitudes differ slightly between the two series, the largest deviations in the spin position are due to the trend differences in the orbital precession. A SPICE kernel for the case of entirely solid satellites has been generated for the $\alpha,\delta$ and $W$ angles, using both the JUP and NOE ephemerides. These 2 kernels are currently available on lara.oma.be/GalileanSat/. Another localization on the SPICE kernel architecture will be available soon. If Europa, Ganymede, or Callisto possess a liquid global ocean beneath the icy shell, certain amplitudes may be amplified, as summarized in Table 18.
@@ Acknowledgments
We thank Sébastien Le Maistre, Tim Van Hoolst, Valery Lainey, Jean-Luc Margot, and Alexander Stark for their helpful comments and discussions. We thank Alfonso Caldiero for testing the SPICE kernel and 2 anonymous reviewers for their valuable suggestions.